Three-quarters of a certain number exceeds two-thirds of it by 3.Find the number

Answer by stanbon(75887)

Three-quarters of a certain number exceeds two-thirds of it by 3.Find the number
----------------------
Let the number be "x":
Equation:
(3/4)x = (2/3)x + 3
---
[(3/4)-(2/3)]x = 3
(1/12)x = 3
x = 36
-------------
Cheers,
Stan H.